OpenStudy (anonymous):

perform the indicated operation and simplify. {[a(x-2y)^3]-[6(x-2y)^4]}/(ab(2y-x)^2)

OpenStudy (anonymous):

\(\huge \frac{a(x-2y)^2-6(x-2y)^4}{ab(2y-x)^2}\\\huge =\frac{(x-2y)^2[a-6(x-2y)^2]}{ab(x-2y)^2}\\\huge =\frac{a-6(x-2y)^2}{ab}\) I use: \(\Large \color{blue}{(a-b)^2=(b-a)^2}\) Thus, \(\Large (2y-x)^2=(x-2y)^2\)